OpenStudy (anonymous):

Find the values of x for (x+3)(x-1)/(x+2)(x-2)>and equal to 0

OpenStudy (anonymous):

\[\frac{ \left( x+3 \right)\left( x-1 \right) }{ \left( x+2 \right)\left( x-2 \right) }\ge 0\] \[ clearly ,x \neq -2,x \neq 2\] \[\frac{ x ^{2}+3x-x-3 }{x ^{2}+2x-2x-4 }=\frac{ x ^{2}+2x-3 }{ x ^{2}-4 }\] \[\frac{ x ^{2}+2x+1-1-3 }{ x ^{2} -4}=\frac{ \left( x+1 \right)^{2}-2^{2} }{ x ^{2}-2^{2} }\] either numerator and denominator both are positive or both are negative case1, both are positive \[\left( x+1 \right)^{2}-2^{2}\ge0,\left| x+1 \right|\ge2\] \[x+1\le-2,x \le -2-1,x \le -3\] \[and x+1\ge 2,x \ge 2-1,x \ge 1 but x \neq 2 x \in [1,2)\cup (2,\infty) \] \[\left| x \right|\le-2,and x \ge2\] \[combining x \in (-\infty,-3]\cup(2,\infty )...(1)\] case2 \[\left( x+1 \right)^{2}-2^{2}<0,\left| x+1 \right|<2,-2<x+1 <2,-3<x <1,but x \neq-2\] \[x \in (-3,-2)\cup (-2,1)\] \[x ^{2}-2^{2}<0,\left| x \right|<2,-2<x <2\] \[combining, x \in (-2,1)... (2)\] (1) and (2) are the solutions. if you find any difficulty in understanding please feel free to ask.